In a factory, three machines produce noise with intensity levels of 85 dB, 90 dB, and 93 dB. When all three are running, what is the intensity level?


1
Expert's answer
2021-05-07T08:11:56-0400

Intensity of sound "\Iota= \Iota_010^{\frac{\beta}{10d\beta}}"


"\Iota_{total} = \Iota_1 + \Iota_2+ \Iota_3"


"\beta_{total}= 10d\beta\log_{10}""({\frac{I_{total}}{I_{0}}})"


"\beta_{total}=" "10d\beta\log_{10}" "(10^{\frac{85dB}{10}} + 10^{\frac{90dB}{10}} +10^{\frac{93dB}{10}})"


"\beta_{total}= 95.2dB"